# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v2

@@ 3.2 Theoretical Analysis
$\bm{\theta}^{\star}$ $={\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\bf w}+\left(\mathbf{I}-{\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}\right)(\widehat{\bm{\theta}}-{\bf z}_{0}),$
@@
$\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})$ $={\bf f}(\widehat{\bm{\theta}},{\bf x})+{{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}\left({\bf w}-{\mathbf{J}}_{\mathcal{X}}\widehat{\bm{\theta}}\right)+\left({{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}-{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x})\right){\bf z}_{0},$ (6)
@@
where ${{\mathbf{K}}}_{{\bf x},\mathcal{X}}\triangleq{{\mathbf{K}}}_{\widehat{\bm{\theta}}}({\bf x},\mathcal{X})\in^{c\times nc}$. Taking ${\bf z}_{0}$ to be a random variable, we form an ensemble of predictors $\{\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\}_{{\bf z}}$, where each $\bm{\theta}^{\star}$ is formed from the projection of a random $\bf z$} onto $\text{Null}({\mathbf{J}}_{\mathcal{X}})$. We require $\bf z$}’s distribution to be symmetric, isotropic, and centered at $\widehat{\bm{\theta}}$, as we do not know a priori which directions contain more information. We take the maximum entropy distribution for a given mean and variance, which is Gaussian [^1] Heavier-tailed distributions matching the above criteria, e.g. logistic distributions, may improve results.. Hence, we let ${\bf z}_{0}\sim\mathcal{N}({\bf 0},\gamma^{2}\mathbf{I})$, for some hyper-parameter $\gamma\in$. The expectation and variance of ˜6, and the distribution of the predictor ${\bf f}(\bm{\theta}^{\star},{\bf x})$, are then
@@
${\mathbb{E}}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\big)=\mathbf{\bm{\mu}}(\widehat{\bm{\theta}},{\bf x})$ $={{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}(\bm{\theta}^{\ddagger}-\widehat{\bm{\theta}})+{\bf f}(\widehat{\bm{\theta}},{\bf x}),$ (7)
@@
$\text{Var}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x})\big)=\mathbf{\bm{\sigma}}^{2}(\widehat{\bm{\theta}},{\bf x})$ $=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}.$ (8)
@@
${\bf f}(\bm{\theta}^{\star},{\bf x})\overset{\text{approx}}{\sim}$ $\mathcal{N}\big(\bm{\mu}(\widehat{\bm{\theta}},{\bf x}),\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})\big).$ (9)
@@
By the full-rank assumption on the Jacobian, this amounts to the conditional distribution of the following normal distribution, conditioned on interpolation ${\bf f}(\bm{\theta}^{\star},\mathcal{X})={\bf y}$,
@@
Therefore, $f(\bm{\theta}^{\star},{\bf x})$ follows a GP with an NTK kernel. Conditioning on ${\bf f}(\bm{\theta}^{\star},\mathcal{X})={\bf y}$ is reasonable, since by construction ${\bf f}(\bm{\theta}^{\star},\mathcal{X})\approx\widetilde{{\bf f}}(\bm{\theta}^{\star},\mathcal{X})$, and under the full-rank assumption of ${\mathbf{J}}_{\mathcal{X}}$, we have $\widetilde{{\bf f}}(\bm{\theta}^{\star},\mathcal{X})={\bf y}$.
@@
Beyond quadratic loss, for a general loss function satisfying the assumptions of Lemma˜3.1, a clear GP posterior interpretation like that in ˜10 may not exist. Nevertheless, we can still derive related insights. If $\widehat{\bm{\theta}}$ is an interpolating solution from initial training, which is common for modern NNs, then as long as ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank, solving ˜3 is equivalent to finding $\bm{\theta}\in^{p}$ such that $\widetilde{{\bf f}}(\bm{\theta},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., find $\bm{\theta}\in^{p}$ such that ${\mathbf{J}}_{\mathcal{X}}(\bm{\theta}-\widehat{\bm{\theta}})={\bf 0}$. So we obtain (9) with $\bm{\mu}(\widehat{\bm{\theta}},{\bf x})={\bf f}(\widehat{\bm{\theta}},{\bf x})$ and $\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}$ as the conditional distribution of ˜10, conditioned on the event ${\bf f}(\bm{\theta}^{\star},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., interpolation with $\ell({\bf f}(\bm{\theta}^{\star},{\bf x}_{i}),{\bf y}_{i})=0$ for $i=1,\ldots,n$.
@@ 4.4 Image Classification - Uncertainty
Figure 3: Violin plot of VMSP, for correctly predicted ID test points, incorrectly predicted ID test points, and OoD test points. Median is shown, with violin width depicting density. Low variance is expected for ID correct points, and large variance for ID incorrect and OoD points.
@@
Figure˜3 presents a violin plot of the VMSP for three test-groups: correctly predicted in-distribution (FashionMNIST, CIFAR-10) test points, incorrectly predicted in-distribution test points, and out-of distribution (MNIST, CIFAR-100) test points. We would expect that there should be, on average, much smaller uncertainty for ID test points that have been correctly predicted, and larger uncertainty for incorrectly predicted ID and OoD test points. We compare against a completely randomized baseline method (BASE), where we sample $10$ standard normal realizations of logits, passed through a softmax. In Table˜10 in Section˜H.8, we display the corresponding median and sample skew values for each method in each test group, to quantify the distribution of VMSP for each test set. Ideally, a method should far outperform the baseline, so we can use the median and sample skew difference between a method and the baseline as a way to compare different methods. We see that NUQLS outperforms all other methods, including the SOTA method DE. In Section˜H.7 we provide additional experimental evaluation of NUQLS. In Figure˜6 we evaluate NUQLS on a ResNet50 trained on both SVHN and ImageNet, displaying the scalability of our method, as well as providing comparison with other competing methods, including Bayesian Deep Ensembles (BDE), Spectral-Normalized Neural Gaussian Process (SNGP), BatchEnsemble (BE), and Stochastic Gradient Langevin Dynamics (SGLD). Against these competing methods, NUQLS performs the strongest.
@@ Appendix A Proofs
    $\bm{\theta}^{\star}$ $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+\sum_{i=1}^{\infty}{\bf v}^{(i)}$
@@
    $\bm{\theta}^{\star}$ $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+\bm{\theta}^{\ddagger}.$
@@
    Let $\bm{\theta}^{\star}$ be any solution to ˜3. The full row-rank assumption on ${\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})$ implies that we must have ${\bf g}(\bm{\theta}^{\star},\mathcal{X})={\bf 0}$, i.e., $\nabla\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x}_{i}),{\bf y}_{i})={\bf 0}$, $i=1,\ldots,n$. By the $\mu$-strong convexity assumption on $\ell(.,{\bf y})$ with respect to its first argument, it is easy to see that $\mathcal{L}(\bm{\theta})$ satisfies the Polyak-Łojasiewicz inequality (Karimi et al., 2016) with constant $2\mu\lambda$ where
@@
    and $\sigma_{\min}({\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i}))$ is the smallest non-zero singular value of ${\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})$. Indeed, let $\bm{\theta}^{\star}$ be any solution to ˜3. From $\mu$-strong convexity of $\ell(.,{\bf y})$ with respect to its first argument, for any $\bm{\theta}$, we have
    $\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})-\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x}_{i}),{\bf y}_{i})$ $\leq\frac{1}{2\mu}\left\|\nabla\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})\right\|^{2}$
@@
    By the smoothness assumption on each $\ell(\widetilde{{\bf f}}(.,{\bf x}_{i}),{\bf y}_{i})$ as well as the interpolating property of $\bm{\theta}^{\star}$, Bassily et al. (2018, Theorem 1) implies that the mini-batch SGD with small enough step size $\eta$ has an exponential convergence rate as

### v3

@@ 3.2 Theoretical Analysis
$\bm{\theta}^{\star}_{{\bf z}}$ $={\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\bf w}+\left(\mathbf{I}-{\mathbf{J}}_{\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}\right)(\widehat{\bm{\theta}}-{\bf z}_{0}),$
@@
$\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})$ $={\bf f}(\widehat{\bm{\theta}},{\bf x})+{{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}\left({\bf w}-{\mathbf{J}}_{\mathcal{X}}\widehat{\bm{\theta}}\right)+\left({{\mathbf{K}}}_{{\bf x},\mathcal{X}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}-{\mathbf{J}}(\widehat{\bm{\theta}},{\bf x})\right){\bf z}_{0},$ (6)
@@
where ${{\mathbf{K}}}_{{\bf x},\mathcal{X}}\triangleq{{\mathbf{K}}}_{\widehat{\bm{\theta}}}({\bf x},\mathcal{X})\in^{c\times nc}$. Taking ${\bf z}_{0}$ to be a random variable, we form an ensemble of predictors $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\}_{{\bf z}}$, where each $\bm{\theta}^{\star}_{{\bf z}}$ is formed from the projection of a random $\bf z$} onto $\text{Null}({\mathbf{J}}_{\mathcal{X}})$. We require $\bf z$}’s distribution to be symmetric, isotropic, and centered at $\widehat{\bm{\theta}}$, as we do not know a priori which directions contain more information. We take the maximum entropy distribution for a given mean and variance, which is Gaussian [^1] Heavier-tailed distributions matching the above criteria, e.g. logistic distributions, may improve results.. Hence, we let ${\bf z}_{0}\sim\mathcal{N}({\bf 0},\gamma^{2}\mathbf{I})$, for some hyper-parameter $\gamma\in$. The expectation and variance of ˜6, and the distribution of the predictor ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})$, are then
@@
${\mathbb{E}}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\big)=\mathbf{\bm{\mu}}(\widehat{\bm{\theta}},{\bf x})$ $={{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{\mathbf{J}}_{\mathcal{X}}(\bm{\theta}^{\ddagger}-\widehat{\bm{\theta}})+{\bf f}(\widehat{\bm{\theta}},{\bf x}),$ (7)
@@
$\text{Var}\big(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\big)=\mathbf{\bm{\sigma}}^{2}(\widehat{\bm{\theta}},{\bf x})$ $=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}.$ (8)
@@
${\bf f}(\bm{\theta}^{\star}_{{\bf z}},{\bf x})\overset{\text{approx}}{\sim}$ $\mathcal{N}\big(\bm{\mu}(\widehat{\bm{\theta}},{\bf x}),\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})\big).$ (9)
@@
By the full-rank assumption on the Jacobian, this amounts to the conditional distribution of the following normal distribution, conditioned on interpolation ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf y}$,
@@
Therefore, $f(\bm{\theta}^{\star}_{{\bf z}},{\bf x})$ follows a GP with an NTK kernel. Conditioning on ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf y}$ is reasonable, since by construction ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})\approx\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})$, and under the full-rank assumption of ${\mathbf{J}}_{\mathcal{X}}$, we have $\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf y}$.
@@
Beyond quadratic loss, for a general loss function satisfying the assumptions of Lemma˜3.1, a clear GP posterior interpretation like that in ˜10 may not exist. Nevertheless, we can still derive related insights. If $\widehat{\bm{\theta}}$ is an interpolating solution from initial training, which is common for modern NNs, then as long as ${\mathbf{J}}_{\mathcal{X}}$ is full row-rank, solving ˜3 is equivalent to finding $\bm{\theta}\in^{p}$ such that $\widetilde{{\bf f}}(\bm{\theta},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., find $\bm{\theta}\in^{p}$ such that ${\mathbf{J}}_{\mathcal{X}}(\bm{\theta}-\widehat{\bm{\theta}})={\bf 0}$. So we obtain (9) with $\bm{\mu}(\widehat{\bm{\theta}},{\bf x})={\bf f}(\widehat{\bm{\theta}},{\bf x})$ and $\bm{\sigma}^{2}(\widehat{\bm{\theta}},{\bf x})=\big({{\mathbf{K}}}_{{\bf x},{\bf x}}-{{\mathbf{K}}}_{{\bf x},\mathcal{X}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}{{\mathbf{K}}}_{\mathcal{X},\mathcal{X}}^{-1}{{\mathbf{K}}}_{{\bf x},\mathcal{X}}\big)\gamma^{2}$ as the conditional distribution of ˜10, conditioned on the event ${\bf f}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf f}(\widehat{\bm{\theta}},\mathcal{X})$, i.e., interpolation with $\ell({\bf f}(\bm{\theta}^{\star}_{{\bf z}},{\bf x}_{i}),{\bf y}_{i})=0$ for $i=1,\ldots,n$.
@@ 4.4 Image Classification - Uncertainty
Figure 3: Violin plot of VMSP, for correctly predicted ID test points, incorrectly predicted ID test points, and OoD test points. Median is shown, with violin width depicting density. Low variance is expected for ID correct points, and large variance for ID incorrect and OoD points. Title of plots gives model and dataset used for training.
@@
Figure˜3 presents a violin plot of the VMSP for three test-groups: correctly predicted in-distribution (FashionMNIST, CIFAR-10, CIFAR-100) test points, incorrectly predicted in-distribution test points, and out-of distribution (MNIST, CIFAR-100, CIFAR-10) test points. We would expect that there should be, on average, much smaller uncertainty for ID test points that have been correctly predicted, and larger uncertainty for incorrectly predicted ID and OoD test points. We compare against a completely randomized baseline method (BASE), where we sample $10$ standard normal realizations of logits, passed through a softmax. In Table˜10 in Section˜H.8, we display the corresponding median and sample skew values for each method in each test group, to quantify the distribution of VMSP for each test set. Ideally, a method should far outperform the baseline, so we can use the median and sample skew difference between a method and the baseline as a way to compare different methods. We see that NUQLS outperforms all other methods, including the SOTA method DE. In Section˜H.7 we provide additional experimental evaluation of NUQLS. In Figure˜6 we evaluate NUQLS on a ResNet50 trained on both SVHN and ImageNet, displaying the scalability of our method, as well as providing comparison with other competing methods, including Bayesian Deep Ensembles (BDE), Spectral-Normalized Neural Gaussian Process (SNGP), BatchEnsemble (BE), and Stochastic Gradient Langevin Dynamics (SGLD). Against these competing methods, NUQLS performs the strongest.
@@ Appendix A Proofs
    $\bm{\theta}^{\star}_{{\bf z}}$ $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+\sum_{i=1}^{\infty}{\bf v}^{(i)}$
@@
    $\bm{\theta}^{\star}_{{\bf z}}$ $=\left(\mathbf{I}-{\mathbf{J}}^{\dagger}(\widehat{\bm{\theta}},\mathcal{X}){\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right){\bf z}+\bm{\theta}^{\ddagger}.$
@@
    Let $\bm{\theta}^{\star}_{{\bf z}}$ be any solution to ˜3. The full row-rank assumption on ${\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})$ implies that we must have ${\bf g}(\bm{\theta}^{\star}_{{\bf z}},\mathcal{X})={\bf 0}$, i.e., $\nabla\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x}_{i}),{\bf y}_{i})={\bf 0}$, $i=1,\ldots,n$. By the $\mu$-strong convexity assumption on $\ell(.,{\bf y})$ with respect to its first argument, it is easy to see that $\mathcal{L}(\bm{\theta})$ satisfies the Polyak-Łojasiewicz inequality (Karimi et al., 2016) with constant $2\mu\lambda$ where
@@
    and $\sigma_{\min}({\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i}))$ is the smallest non-zero singular value of ${\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})$. Indeed, let $\bm{\theta}^{\star}_{{\bf z}}$ be any solution to ˜3. From $\mu$-strong convexity of $\ell(.,{\bf y})$ with respect to its first argument, for any $\bm{\theta}$, we have
    $\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})-\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star}_{{\bf z}},{\bf x}_{i}),{\bf y}_{i})$ $\leq\frac{1}{2\mu}\left\|\nabla\ell(\widetilde{{\bf f}}(\bm{\theta},{\bf x}_{i}),{\bf y}_{i})\right\|^{2}$
@@
    By the smoothness assumption on each $\ell(\widetilde{{\bf f}}(.,{\bf x}_{i}),{\bf y}_{i})$ as well as the interpolating property of $\bm{\theta}^{\star}_{{\bf z}}$, Bassily et al. (2018, Theorem 1) implies that the mini-batch SGD with small enough step size $\eta$ has an exponential convergence rate as